A gas mixture that has the following composition in percent by volume: 10.2% CO2,

21.0% CO, 18.5% H2, 2.3% CH4 and 48.0% N2.

(a) What is the density of this gas at 24 °C and 755 mmHg, in grams per liter?

(b) What is the partial pressure of CO in this mixture at STP?

(c) What volume of air is required for the complete combustion of this mixture of gas? (Air

contains 21% O2 by volume)


1
Expert's answer
2021-12-20T11:26:03-0500

The  density of a substance changes with change in temperature. This is because the volume increases as temperature increases and decreases as the temperature decreases. The mass is not affected by temperature or volume. We need to determine the volume at STP in order to determine the density at STP.

According to Dalton's Law of partial pressure.The partial pressure of gas ∝ Mole fraction of gas (X gas)XCO=nCO+nN2.Hence the partial pressure is 0.5 V(H2+CH4+CO+C2H4+non−combustible mixture)=1000lit
Volume % of H2​=50%⇒ volume H2​=500lit
Volume % of CH4​=35%⇒ volume CH4​=350lit
Volume % of CO=8%⇒ volume CO=80lit
Volume % of non-combustible mixture =5%⇒ volume =50litH2+1/2O2​→H2O
500 lit requires 250 lit of O2​
CH4​+2O2​→CO2​+2H2​O350 lit requires 350×2=700 lit of O2​
CO+1/2O2​→CO2​80 lit requires 80×2=160 lit of O2​
C2​H4​+3O2​→2CO2​+2H2​O20 lit requires 60 lit of O2​
∴ Total volume of O2​ (required)=250lit+700lit+160lit+60lit=1170lit
∴ If volume of air =xlitThen x×21%=1050⇒x=21
1170×100​=5571lit=5.57 m3